The second of three numbers is 1 less than the first. The third number is 5 less than the second. If the first number is twice as large as the third, find the three numbers.: The second of three numbers is 1 less than the first. The third number is 5 less than the second. If the first number is twice as large as the third, find the three numbers.
Answer by KnightOwlTutor(216)

#1=X
#2=x-1
#3=x-1-5=x-6
X=2(X-6)
Use the distributive property
x=2x-12
subtract 2x from both sides
-x=-12
x=12
#1=12
#2=11
#3=6
Let's plug in the values to check our answer
12=2(6)
12=12